• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 08:03:21 on June26,2002

    Entry number 86810

    keyword=Shift Summary

    00:00 MCC still working on the machine

    01:00 Q1 Tripped and was reset

    02:15 MCC cannot deliver more than 8 uA. Optics is being called in.

    03:40 Good beam to Hall A (100 uA, fast feedback on, position 1H04B -0.1:0.0
    target OTR looks good) 1H04A is a bit off, but MCC was unable to fix it.

    Kinematics Q2_jkN file 1814 (high prescales) 15cm LD
    file 1815

    4:05 Went to change kinematics and spectrometer bogie control failed.
    Gary arrived, a controlled access was made, and the problem fixed.

    5:40 Started taking physics beam again

    Kinematic Q2_k40 file 1816 Optics Target
    file 1817 15cm LD Target 1M events
    file 1818 15cm LD Target 0.9 M events

    7:55 North Linac just dropped to power permit.